# OmniUMI: Towards Physically Grounded Robot Learning via Human-Aligned Multimodal Interaction

### 12 Apr 2026 v1

@@ Abstract
UMI-style interfaces enable scalable robot learning, but existing systems remain largely visuomotor, relying primarily on RGB observations and trajectory while providing only limited access to physical interaction signals. This becomes a fundamental limitation in contact-rich manipulation, where success depends on contact dynamics such as tactile interaction, internal grasping force, and external interaction wrench that are difficult to infer from vision alone. We present OmniUMI, a unified framework for physically grounded robot learning via human-aligned multimodal interaction. OmniUMI synchronously captures RGB, depth, trajectory, tactile sensing, internal grasping force, and external interaction wrench within a compact handheld system, while maintaining collection–deployment consistency through a shared embodiment design. To support human-aligned demonstration, OmniUMI provides dual-force feedback through bilateral gripper feedback and natural perception of external interaction wrench in the handheld embodiment. Built on this interface, we extend diffusion policy with visual, tactile, and force-related observations, and deploy the learned policy through impedance-based execution for unified regulation of motion and contact behavior. Experiments demonstrate reliable sensing and strong downstream performance on force-sensitive pick-and-place, interactive surface erasing, and tactile-informed selective release. Overall, OmniUMI combines physically grounded multimodal data acquisition with human-aligned interaction, providing a scalable foundation for learning contact-rich manipulation.
Figure 1: OmniUMI overview. Left: a unified multimodal handheld interface for physically grounded data acquisition, synchronously capturing RGB, depth, trajectory, tactile sensing, internal grasping force, and external interaction wrench, with dual-force feedback for grounded human demonstration. Middle: multimodal policy learning over visual, tactile, and force-related observations. Right: impedance-compatible deployment, where policy outputs are translated into virtual targets and executed for unified regulation of motion and contact behavior in contact-rich manipulation.
@@ 1 Introduction
In this work, we present OmniUMI, a unified framework for physically grounded robot learning via human-aligned multimodal interaction. Rather than treating tactile, force, and motion signals as independent add-ons, OmniUMI is designed around the concept of interaction acquisition, in which policies are trained directly on physically meaningful signals of contact and force collected through a human-aligned demonstration interface. To realize this goal, our framework is built around three tightly coupled components: (i) a unified multimodal handheld interface for physically grounded data acquisition, (ii) a human-aligned demonstration pipeline centered on force interaction, and (iii) a multimodal learning and impedance-compatible deployment framework. Together, these components establish a consistent pipeline from multimodal human interaction to contact-rich robot execution.
@@
At the demonstration level, OmniUMI emphasizes human-aligned interaction. To better align recorded demonstrations with operator intent during contact-rich manipulation, OmniUMI provides dual-force feedback: grasping force is rendered through a custom master–slave gripper with bilateral control, while the handheld form factor naturally preserves intuitive perception of external interaction wrench. Together, these mechanisms allow operators to simultaneously feel, modulate, and record both internal grasping force and external interaction wrench. In this sense, OmniUMI does not merely measure physical variables; it organizes their acquisition around a human-aligned interface for natural interaction regulation.
@@
  • A human-aligned interaction acquisition paradigm with dual-force feedback. We introduce a demonstration interface in which internal grasping force is grounded through bilateral force feedback in a custom master–slave gripper, while external interaction wrench is naturally grounded through the handheld embodiment itself, enabling both signals to be perceived, modulated, and recorded during demonstration and improving alignment between recorded interaction signals and human intent.
@@ 2.2 From Visuomotor Interfaces to Physically Grounded Multimodal Interaction
However, prior efforts are still largely modality-centric. Most systems augment existing pipelines with one additional sensing channel at a time, validating the usefulness of that modality in isolation, but without fully addressing how multiple heterogeneous interaction variables should be jointly integrated into a compact, reusable, and deployment-compatible robot-free interface. This leaves open the broader systems question that motivates our work: how to move from scalable visuomotor interfaces to scalable physically grounded robot learning via human-aligned multimodal interaction.
@@ 2.3 Tactile Sensing in Robot-free Manipulation Interfaces
However, the key challenge is no longer the utility of tactile sensing itself, but how to incorporate it into a practical robot-free interface. Many existing designs integrate tactile sensors directly into the gripper fingertips [4], which is mechanically simple but compromises gripper usability, limits small-object manipulation, and exposes sensors to damage. At the same time, tactile signals are often high-dimensional and difficult to integrate into policy learning without careful representation design [8]. Therefore, while prior work establishes the importance of tactile sensing, it does not fully resolve how to embed tactile perception into a robust, reusable, and scalable interaction interface.
@@ 2.6 Summary and Our Position
Our work is positioned beyond modality-specific extensions. Rather than introducing only an additional sensing channel, we aim to build a unified robot-free framework that jointly supports RGB, depth, trajectory, tactile sensing, internal grasping force, and external interaction wrench within a single reusable system. Moreover, we emphasize both physically grounded multimodal data acquisition and human-aligned interaction acquisition: the former by explicitly capturing physically meaningful interaction variables, and the latter by organizing their collection around natural human perception and regulation through dual-force feedback. Combined with controller-compatible multimodal policy learning and impedance-based deployment, our system is not merely a tactile- or force-augmented variant of existing robot-free interfaces, but a scalable framework for physically grounded robot learning via human-aligned multimodal interaction.
@@ 3.1 Overview
Our approach is built around three tightly coupled components: (i) a unified multimodal handheld interface for physically grounded data acquisition, (ii) a human-aligned interaction acquisition pipeline centered on dual-force feedback, and (iii) a multimodal learning and impedance-compatible deployment framework. Together, these components establish a consistent pipeline from robot-free multimodal human interaction to contact-rich robot execution.
Rather than treating tactile, force, and motion signals as independent add-ons, OmniUMI is designed around the concept of interaction acquisition, where policies are trained directly on physically meaningful signals of contact and force. This perspective shapes not only the sensing embodiment, but also the structure of the collected data, the semantics of the learned policy interface, and the way learned actions are executed on the robot.
@@ 3.2 Design Objectives
Human-Aligned Interaction Acquisition. Collecting more sensing channels is not sufficient if the resulting data do not faithfully reflect how humans regulate interaction. OmniUMI therefore aims to acquire physical variables through an interface that allows the operator to naturally perceive and modulate them during demonstration, especially for grasp regulation and contact-rich manipulation. This objective motivates our emphasis on dual-force feedback, through which internal grasping force is rendered by bilateral gripper interaction while external interaction wrench remains naturally perceivable through the handheld embodiment.
@@ 3.4 Physically Grounded and Human-Aligned Interaction Acquisition
With the system embodiment established, the next step is to acquire data that are both physically grounded and human-aligned. In OmniUMI, this means that recorded multimodal signals should capture physically meaningful robot–environment interaction variables, while remaining consistent with how the operator perceives and regulates contact during demonstration. This requirement is especially important in contact-rich manipulation, where task success depends not only on motion geometry, but also on how grasp force, interaction wrench, and local contact evolve throughout the manipulation process. For this reason, OmniUMI organizes data acquisition around internal force and external force as the primary grounded interaction variables, while treating tactile, RGB, depth, and trajectory as complementary supporting observations.
@@
$\tau_{s}=K_{p}\left(b\,\theta_{m}-\theta_{s}\right)+K_{d}\left(b\,\dot{\theta}_{m}-\dot{\theta}_{s}\right)$ (2)
$\tau_{m}=-\frac{1}{a}\left[K_{p}\left(b\,\theta_{m}-\theta_{s}\right)+K_{d}\left(b\,\dot{\theta}_{m}-\dot{\theta}_{s}\right)\right]$ (3)
where $\theta_{m},\theta_{s}$ and $\dot{\theta}_{m},\dot{\theta}_{s}$ denote the motor position and velocity of the master and slave grippers, respectively, $K_{p}$ and $K_{d}$ are feedback gains, $b$ is a position scaling factor, and $a$ controls the reflected torque magnitude. Through this closed interaction loop, the operator can feel contact onset, object resistance, and grasp transitions directly at the master side rather than inferring them only from visual cues.
This bilateral design improves more than user experience. It directly improves the quality of the collected data by making the internal-force signal not only measurable, but also human-aligned. In this sense, internal force in OmniUMI is not merely sensed; it is acquired through an interaction loop in which the human can both generate and regulate the variable being recorded. At the same time, because the gripper is motorized rather than mechanically triggered, the design also avoids introducing actuation-induced disturbances into the external force/torque sensor. As a result, the same hardware choice supports both physically grounded internal-force acquisition and human-aligned interaction during demonstration.
More generally, this design illustrates that interaction variables become human-aligned not only when they are sensed, but when they are embedded in a feedback loop through which operators can directly regulate them.
@@
Tactile and Visual Observations.
In addition to dual-force interaction acquisition, OmniUMI also acquires tactile images together with RGB, depth, and trajectory observations as supporting modalities. These channels provide complementary information: tactile sensing captures local contact appearance, RGB and depth provide scene-level context and geometric structure, and trajectory provides motion reference. From a system perspective, their role is not to replace force interaction variables, but to complement them. Their synchronized combination with internal and external force allows the collected dataset to retain both global context and local interaction detail within one robot-free multimodal interface.
Since the main contribution of this work does not lie in advancing these sensing channels individually, we keep this component lightweight and focus instead on their synchronized integration into the overall multimodal interface. In the current system, tactile observations are directly used as image-based inputs to the policy, without introducing an additional tactile-specific representation module. This choice keeps the learning pipeline simple and allows the main methodological emphasis to remain on physically grounded multimodal data acquisition, human-aligned interaction acquisition, and deployment-compatible control.
@@ 3.5 Multimodal Policy Learning
As illustrated in Fig. 3, the policy conditions the diffusion model on both multimodal observations and diffusion-step embeddings, and predicts action chunks that are executed in a receding-horizon manner. The policy maps synchronized multimodal observations to action predictions:
@@
$\mathbf{o}_{t}=\left\{I^{rgb}_{t},\,I^{depth}_{t},\,I^{tac}_{t},\,\mathbf{p}_{t},\,\mathbf{f}^{int}_{t},\,\mathbf{f}^{ext}_{t}\right\}$ (6)
where the observation includes synchronized visual, tactile, kinematic, and force-related signals.
@@
$\mathbf{a}_{t}=\left[\Delta x_{t},\,\Delta y_{t},\,\Delta z_{t},\,\mathbf{r}^{6D}_{t},\,w^{g}_{t}\right]$ (7)
consisting of end-effector translation in xyzxyz, a 6D rotation representation, and a virtual gripper width.
@@ 4 Experiments
We evaluate OmniUMI from three complementary perspectives: (1) whether the proposed interface provides reliable physically grounded sensing for contact-rich interaction, (2) whether the proposed human-aligned interaction design improves the quality of collected demonstrations, and (3) whether the resulting multimodal data improve downstream policy learning on representative contact-rich manipulation tasks. Consistent with the central claim of this work, the downstream evaluations are organized around three key interaction capabilities enabled by the proposed framework: grasp-force-aware manipulation, wrench-informed surface interaction, and tactile-informed fine-grained control. Taken together, these experiments assess OmniUMI not merely as a multimodal sensing platform, but as a unified robot-free framework for physically grounded robot learning via human-aligned multimodal interaction.
@@ 4.2 Human-Aligned Demonstration Quality
We next evaluate whether OmniUMI improves the quality of collected demonstrations through its human-aligned interaction design. We consider two complementary aspects: comparison with conventional teleoperation in a contact-rich wiping task, and comparison of grasp-force collection with and without bilateral force feedback. The first evaluation focuses on whether the handheld multimodal interface yields more natural contact regulation during demonstration, while the second directly tests whether dual-force feedback improves the quality of recorded internal-force trajectories.
Comparison with teleoperation-based wiping data collection.
To assess the practical advantages of OmniUMI over conventional teleoperation-based data collection, we compare three data acquisition settings for the whiteboard wiping task: direct human demonstration, teleoperation, and OmniUMI. For each setting, we collect several wiping trajectories and analyze both the collection time and the recorded force signals.
Because whiteboard wiping is a contact-dominant task, the most relevant force component is the normal contact force along the $z$ axis. We therefore focus on the recorded $F_{z}$ profiles as the primary indicator of interaction quality.
Figure 5: Representative $F_{z}$ profiles collected during whiteboard wiping. We compare direct human demonstration, teleoperation, and OmniUMI. Since whiteboard wiping is dominated by normal contact force, the $z$-axis force provides the most informative comparison. OmniUMI produces force profiles and task durations that are much closer to direct human demonstrations, whereas teleoperation requires substantially longer collection time and exhibits severe force oscillations.
As shown in Fig. 5, the wiping trajectories collected with OmniUMI exhibit average task durations and $F_{z}$ profiles that are much closer to those of direct human demonstrations. In contrast, the teleoperation baseline requires noticeably longer collection time and produces highly oscillatory force signals, indicating unstable contact regulation during wiping.
These observations suggest that, without direct and physically grounded interaction feedback, teleoperation users rely mainly on visual cues and tend to overcompensate at contact, which leads to inefficient demonstrations and severe force fluctuations. By contrast, OmniUMI enables more natural contact modulation and produces smoother, more human-like interaction data. This result supports the claim that the handheld embodiment and interaction-centered sensing design improve human alignment during demonstration.
Comparison of grasp-force collection with and without bilateral feedback.
To further evaluate the role of bilateral force feedback in human-aligned interaction acquisition, we compare three demonstration settings for grasp-force collection: direct human demonstration, OmniUMI without bilateral force feedback, and the full OmniUMI system with bilateral force feedback. For each setting, we collect several grasping trajectories and compare the recorded internal grasp-force profiles.
Because the target interaction variable in grasping is internal force, we focus on the temporal evolution of the recorded grasp-force signal as the primary indicator of demonstration quality.
@@
Figure 6: Representative internal grasp-force profiles collected under three demonstration settings: direct human demonstration, OmniUMI without bilateral force feedback, and OmniUMI with bilateral force feedback. The version with bilateral feedback produces grasp-force trajectories that are substantially closer to direct human demonstrations, while removing bilateral feedback leads to less stable and less human-like force modulation.
@@
As shown in Fig. 6, the grasp-force trajectories collected with the full OmniUMI system are much closer to those obtained from direct human demonstrations, both in overall magnitude and in temporal modulation patterns. In contrast, when bilateral force feedback is removed, the collected force trajectories become less stable and exhibit larger deviations from the human reference, indicating poorer grasp-force regulation during demonstration.
These observations suggest that bilateral force feedback plays a critical role in making the interaction acquisition process human-aligned. Without such feedback, the operator can command gripper motion but cannot directly feel or regulate the evolving grasp interaction, which leads to less natural force modulation. By contrast, the full OmniUMI system enables the operator to directly perceive and modulate grasp force during demonstration, resulting in smoother and more human-like internal-force trajectories. This makes bilateral feedback essential for collecting high-quality force-aware demonstrations for contact-rich manipulation.

### 22 Apr 2026 v2

@@ Abstract
UMI-style interfaces enable scalable robot learning, but existing systems remain largely visuomotor, relying primarily on RGB observations and trajectory while providing only limited access to physical interaction signals. This becomes a fundamental limitation in contact-rich manipulation, where success depends on contact dynamics such as tactile interaction, internal grasping force, and external interaction wrench that are difficult to infer from vision alone. We present OmniUMI, a unified framework for physically grounded robot learning via human-aligned multimodal interaction. [^1] Project page: https://baai-aether.github.io/OmniUMI OmniUMI synchronously captures RGB, depth, trajectory, tactile sensing, internal grasping force, and external interaction wrench within a compact handheld system, while maintaining collection–deployment consistency through a shared embodiment design. To support human-aligned demonstration, OmniUMI enables natural perception and modulation of internal grasping force, external interaction wrench, and tactile interaction through bilateral gripper feedback and the handheld embodiment. Built on this interface, we extend diffusion policy with visual, tactile, and force-related observations, and deploy the learned policy through impedance-based execution for unified regulation of motion and contact behavior. Experiments demonstrate reliable sensing and strong downstream performance on force-sensitive pick-and-place, interactive surface erasing, and tactile-informed selective release. Overall, OmniUMI combines physically grounded multimodal data acquisition with human-aligned interaction, providing a scalable foundation for learning contact-rich manipulation.
Figure 1: OmniUMI overview. Left: a unified multimodal handheld interface for physically grounded data acquisition, synchronously capturing RGB, depth, trajectory, tactile sensing, internal grasping force, and external interaction wrench, with bilateral gripper feedback and a handheld embodiment for human-aligned demonstration. Middle: multimodal policy learning over visual, tactile, and force-related observations. Right: impedance-compatible deployment, where policy outputs are translated into virtual targets and executed for unified regulation of motion and contact behavior in contact-rich manipulation.
@@ 1 Introduction
In this work, we present OmniUMI, a unified framework for physically grounded robot learning via human-aligned multimodal interaction. Rather than treating tactile, force, and motion signals as independent add-ons, OmniUMI is designed around the concept of interaction acquisition, in which policies are trained directly on physically meaningful multimodal interaction signals collected through a human-aligned demonstration interface. To realize this goal, our framework is built around three tightly coupled components: (i) a unified multimodal handheld interface for physically grounded data acquisition, (ii) a human-aligned demonstration pipeline centered on multimodal physical interaction, and (iii) a multimodal learning and impedance-compatible deployment framework. Together, these components establish a consistent pipeline from multimodal human interaction to contact-rich robot execution.
@@
At the demonstration level, OmniUMI emphasizes human-aligned interaction. To better align recorded demonstrations with operator intent during contact-rich manipulation, OmniUMI grounds multimodal physical interaction through bilateral gripper feedback and the handheld embodiment: internal grasping force is rendered through a custom master–slave gripper with bilateral control, while the handheld form factor naturally preserves perception of external interaction wrench and tactile interaction. Together, these mechanisms allow operators to naturally perceive, modulate, and record internal grasping force, external interaction wrench, and tactile interaction during demonstration. In this sense, OmniUMI does not merely measure physical variables; it organizes their acquisition around a human-aligned interface for natural interaction regulation.
@@
  • A human-aligned multimodal interaction acquisition paradigm via bilateral gripper feedback and handheld embodiment. We introduce a demonstration interface in which multimodal physical interaction is grounded through bilateral gripper feedback and the handheld embodiment: internal grasping force is rendered through bilateral control in a custom master–slave gripper, while external interaction wrench and tactile interaction are naturally grounded through the handheld embodiment, enabling these interaction signals to be perceived, modulated, and recorded during demonstration and improving alignment between recorded interaction signals and human intent.
@@ 2.2 From Visuomotor Interfaces to Physically Grounded Multimodal Interaction
However, prior efforts are still largely modality-centric. Most systems augment existing pipelines with one additional sensing channel at a time, validating the usefulness of that modality in isolation, but without fully addressing how multiple heterogeneous interaction variables should be jointly integrated into a compact, reusable, and deployment-compatible robot-free interface. This leaves open the broader systems question that motivates our work: how to move from scalable visuomotor interfaces to scalable physically grounded robot learning via human-aligned multimodal interaction, in which tactile interaction, internal grasping force, and external interaction wrench are not only measured, but also naturally perceived and regulated during demonstration.
@@ 2.3 Tactile Sensing in Robot-free Manipulation Interfaces
However, the key challenge is no longer the utility of tactile sensing itself, but how to incorporate it into a practical robot-free interface. Many existing designs integrate tactile sensors directly into the gripper fingertips [4], which is mechanically simple but compromises gripper usability, limits small-object manipulation, and exposes sensors to damage. At the same time, tactile signals are often high-dimensional and difficult to integrate into policy learning without careful representation design [8]. Therefore, while prior work establishes the importance of tactile sensing, it does not fully resolve how to embed tactile perception into a robust, reusable, and scalable interaction interface that also preserves natural and human-aligned tactile interaction during demonstration.
@@ 2.6 Summary and Our Position
Our work is positioned beyond modality-specific extensions. Rather than introducing only an additional sensing channel, we aim to build a unified robot-free framework that jointly supports RGB, depth, trajectory, tactile sensing, internal grasping force, and external interaction wrench within a single reusable system. Moreover, we emphasize both physically grounded multimodal data acquisition and human-aligned interaction acquisition: the former by explicitly capturing physically meaningful interaction variables, and the latter by organizing their collection around natural human perception and regulation of internal grasping force, external interaction wrench, and tactile interaction through bilateral gripper feedback and the handheld embodiment. Combined with controller-compatible multimodal policy learning and impedance-based deployment, our system is not merely a tactile- or force-augmented variant of existing robot-free interfaces, but a scalable framework for physically grounded robot learning via human-aligned multimodal interaction.
@@ 3.1 Overview
Our approach is built around three tightly coupled components: (i) a unified multimodal handheld interface for physically grounded data acquisition, (ii) a human-aligned multimodal interaction acquisition pipeline, and (iii) a multimodal learning and impedance-compatible deployment framework. Together, these components establish a consistent pipeline from robot-free multimodal human interaction to contact-rich robot execution.
Rather than treating tactile, force, and motion signals as independent add-ons, OmniUMI is designed around the concept of interaction acquisition, where policies are trained directly on physically meaningful multimodal interaction signals. This perspective shapes not only the sensing embodiment, but also the structure of the collected data, the semantics of the learned policy interface, and the way learned actions are executed on the robot.
@@ 3.2 Design Objectives
Human-Aligned Interaction Acquisition. Collecting more sensing channels is not sufficient if the resulting data do not faithfully reflect how humans regulate interaction. OmniUMI therefore aims to acquire physical variables through an interface that allows the operator to naturally perceive and modulate internal grasping force, external interaction wrench, and tactile interaction during demonstration. This objective motivates our co-design of bilateral gripper feedback and the handheld embodiment, through which these interaction variables remain aligned with the operator’s natural contact experience during data collection.
@@ 3.4 Physically Grounded and Human-Aligned Interaction Acquisition
With the system embodiment established, the next step is to acquire data that are both physically grounded and human-aligned. In OmniUMI, this means that recorded multimodal signals should capture physically meaningful robot–environment interaction variables, while remaining consistent with how the operator perceives and regulates contact during demonstration. This requirement is especially important in contact-rich manipulation, where task success depends not only on motion geometry, but also on how grasp force, interaction wrench, and local contact evolve throughout the manipulation process. For this reason, OmniUMI organizes data acquisition around three human-aligned physical interaction channels—internal grasping force, external interaction wrench, and tactile interaction—while using RGB, depth, and trajectory as complementary supporting observations.
@@
$\tau_{s}=K_{p}\left(b\,\theta_{m}-\theta_{s}-\delta\right)+K_{d}\left(b\,\dot{\theta}_{m}-\dot{\theta}_{s}\right)$ (2)
$\tau_{m}=-\frac{1}{a}\,\tilde{\tau}_{s}+B_{l}\,\dot{\theta}_{m}$ (3)
where $\theta_{m},\theta_{s}$ and $\dot{\theta}_{m},\dot{\theta}_{s}$ denote the motor position and velocity of the master and slave grippers, respectively, $\delta$ is the installation offset between the master and slave sides, $K_{p}$ and $K_{d}$ are feedback gains, $b$ is a position scaling factor, $\tilde{\tau}_{s}$ is the filtered measured torque at the slave side, $a$ controls the reflected torque magnitude, and $B_{l}$ is a local velocity-feedback term on the master side. Under the present sign convention, this term acts not as additional damping injection, but as a compensation of the effective resistance felt at the master side. In practice, it offsets part of the sluggishness introduced by torque reflection, thereby reducing the viscous drag perceived by the operator and making master-side manipulation lighter and smoother. Through this closed interaction loop, the operator can feel contact onset, object resistance, and grasp transitions directly at the master side rather than inferring them only from visual cues.
This bilateral design improves more than user experience. It directly improves the quality of the collected data by making the internal-force signal not only measurable, but also human-aligned. In this sense, internal force in OmniUMI is not merely sensed; it is acquired through an interaction loop in which the human can both generate and regulate the variable being recorded. More specifically, the slave side realizes PD-based position tracking of the master-side grasp motion, while the master side renders the reflected torque induced by contact at the slave side, allowing the operator to directly perceive grasp interaction during demonstration. The local velocity-feedback term further improves operability by compensating part of the reflected impedance, rather than by introducing extra damping. At the same time, because the gripper is motorized rather than mechanically triggered, the design also avoids introducing actuation-induced disturbances into the external force/torque sensor. As a result, the same hardware choice supports both physically grounded internal-force acquisition and human-aligned interaction during demonstration.
More generally, this design illustrates that interaction variables become human-aligned not only when they are sensed, but when the embodiment allows operators to naturally perceive and regulate them during demonstration.
@@
Tactile Interaction and Visual Observations.
In addition to internal grasping force and external interaction wrench, OmniUMI also acquires tactile interaction together with RGB, depth, and trajectory observations. Tactile sensing is not treated merely as an auxiliary modality, but as a human-aligned record of local contact deformation during manipulation. These channels provide complementary information: tactile sensing captures local contact deformation and contact evolution, RGB and depth provide scene-level context and geometric structure, and trajectory provides motion reference. From a system perspective, tactile interaction complements internal and external force by preserving local contact detail, while RGB, depth, and trajectory provide supporting context for multimodal policy learning. Their synchronized combination with internal and external force allows the collected dataset to retain both global context and local interaction detail within one robot-free multimodal interface.
Since the main contribution of this work does not lie in advancing these sensing channels individually, we keep this component lightweight and focus instead on their synchronized integration into the overall multimodal interface. In the current system, tactile images are first encoded into low-dimensional tactile embeddings, which are then combined with other low-dimensional observations before being fed to the policy. This choice preserves a compact policy interface while allowing the main methodological emphasis to remain on physically grounded multimodal data acquisition, human-aligned interaction acquisition, and deployment-compatible control.
@@ 3.5 Multimodal Policy Learning
As illustrated in Fig. 3, the policy conditions the diffusion model on both multimodal observations and diffusion-step embeddings, and predicts action chunks that are executed in a receding-horizon manner. We first encode the tactile image into a compact tactile embedding
$\mathbf{z}^{tac}_{t}=\phi_{tac}\!\left(I^{tac}_{t}\right),$ (5)
where $\phi_{tac}$ denotes the tactile encoder.
The policy then maps synchronized multimodal observations to action predictions:
@@
$\mathbf{o}_{t}=\left\{I^{rgb}_{t},\,I^{depth}_{t},\,\mathbf{z}^{tac}_{t},\,\mathbf{p}_{t},\,\mathbf{f}^{int}_{t},\,\mathbf{f}^{ext}_{t}\right\}$ (7)
where the observation includes synchronized visual inputs, tactile embeddings, kinematic signals, and force-related signals.
@@
$\mathbf{a}_{t}=\left[\Delta x_{t},\,\Delta y_{t},\,\Delta z_{t},\,\mathbf{r}^{6D}_{t},\,f^{x}_{t},\,f^{y}_{t},\,f^{z}_{t},\,w^{g}_{t}\right]$ (8)
consisting of end-effector translation in xyzxyz, a 6D rotation representation, a predicted Cartesian interaction force $\mathbf{f}_{t}=[f^{x}_{t},f^{y}_{t},f^{z}_{t}]^{\top}$, and a virtual gripper width.
@@ 4 Experiments
We evaluate OmniUMI from three complementary perspectives: (1) whether the proposed interface provides reliable physically grounded sensing for contact-rich interaction, (2) whether the proposed human-aligned interaction design improves the quality of collected demonstrations across internal grasp-force, external-force, and tactile alignment, and (3) whether the resulting multimodal data improve downstream policy learning on representative contact-rich manipulation tasks. Consistent with the central claim of this work, the downstream evaluations are organized around three key interaction capabilities enabled by the proposed framework: grasp-force-aware manipulation, wrench-informed surface interaction, and tactile-informed fine-grained control. Taken together, these experiments assess OmniUMI not merely as a multimodal sensing platform, but as a unified robot-free framework for physically grounded robot learning via human-aligned multimodal interaction.
@@ 4.2 Human-Aligned Demonstration Quality
We next evaluate whether OmniUMI improves the quality of collected demonstrations through its human-aligned interaction design. Rather than using different tasks for different modalities, we use a single contact-rich manipulation task to examine three forms of alignment in a unified manner: internal grasp-force alignment, external-force alignment, and tactile alignment. The task requires the operator to pick up a whiteboard eraser, erase red marker traces on the whiteboard, and then place the eraser back. Across all three evaluations, we compare the same three demonstration settings: direct human demonstration, teleoperation without bilateral force feedback, and OmniUMI with bilateral gripper feedback and a handheld embodiment. For each setting, we analyze one representative trajectory from the same task and examine whether the recorded interaction signals are closer to natural human behavior across multiple physical channels, including internal grasp force, external contact force, and tactile deformation.
Internal grasp-force alignment.
We first evaluate human alignment from the perspective of internal grasp force. Although the overall task includes grasping, wiping, and object placement, the internal grasp-force signal is most directly related to how the operator picks up, holds, and releases the eraser throughout the task. We therefore focus on the temporal evolution of the recorded grasp-force trajectory as the primary indicator of internal-force alignment.
@@
Figure 5: Representative internal grasp-force trajectories recorded during the whiteboard eraser pick–wipe–place task under three demonstration settings: direct human demonstration, teleoperation without bilateral force feedback, and OmniUMI with bilateral gripper feedback and a handheld embodiment. OmniUMI produces grasp-force trajectories that are substantially closer to direct human demonstrations in both overall magnitude and temporal modulation, whereas teleoperation without bilateral force feedback leads to less stable and less human-like force regulation.
@@
As shown in Fig. 5, the internal grasp-force trajectory collected with OmniUMI more closely matches that of direct human demonstration, both in overall magnitude and in temporal variation. In contrast, the teleoperation baseline produces a less stable force trajectory and larger deviations from the human reference, indicating poorer regulation of grasp interaction during demonstration.
This result suggests that internal grasp-force alignment cannot be achieved reliably through visual observation alone. Without bilateral force feedback, teleoperation users must rely primarily on visual observation and therefore have difficulty reproducing the timing and magnitude of human-like grasp-force modulation. By contrast, OmniUMI combines bilateral gripper feedback with a handheld embodiment, enabling the operator to regulate grasp interaction in a more natural and embodied manner and resulting in smoother and more human-like internal-force trajectories.
External-force alignment.
We next evaluate human alignment from the perspective of external contact force using the same task. Among the different phases of the trajectory, the wiping stage is the most contact-critical, because successful erasing depends on maintaining stable contact with the whiteboard surface while removing the red marker traces. Accordingly, we focus on the recorded normal contact force along the $z$ axis during wiping, i.e., the $F_{z}$ profile, as the primary indicator of external-force alignment.
Figure 6: Representative $F_{z}$ profiles recorded during the wiping phase of the whiteboard eraser pick–wipe–place task under three demonstration settings: direct human demonstration, teleoperation without bilateral force feedback, and OmniUMI with bilateral gripper feedback and a handheld embodiment. Since effective whiteboard cleaning depends critically on stable normal contact during erasing, the $z$-axis force provides the most informative comparison. OmniUMI produces force trajectories that are much closer to direct human demonstrations, whereas teleoperation exhibits more severe force oscillations during wiping.
As shown in Fig. 6, the $F_{z}$ trajectory collected with OmniUMI during the wiping phase is much closer to that of direct human demonstration. In contrast, the teleoperation baseline produces highly oscillatory force signals, indicating unstable contact regulation in the most interaction-sensitive stage of the task.
These observations suggest that, without bilateral force feedback, teleoperation users rely mainly on visual cues and tend to overcompensate at contact, which leads to unstable force modulation. By contrast, OmniUMI combines bilateral gripper feedback with a handheld structure that preserves more natural embodied contact regulation during wiping, thereby producing external-force trajectories that better match direct human demonstrations.
Tactile alignment.
We further evaluate human alignment from the perspective of tactile interaction, again using the same whiteboard eraser pick–wipe–place task. To quantify tactile interaction, we use the marker-motion magnitude on the tactile image as the evaluation metric. Specifically, let $\Delta\mathbf{m}_{t}\in\mathbb{R}^{126}$ denote the marker offset vector at time step $t$, obtained by concatenating the 2D displacements of all tactile markers. We then compute its $\ell_{2}$ norm, $\|\Delta\mathbf{m}_{t}\|_{2}$, which measures the overall deformation magnitude of the tactile image over time. Because this quantity directly reflects the strength and temporal variation of local contact-induced deformation, it provides an informative proxy for tactile alignment.
Figure 7: Representative tactile interaction trajectories recorded during the whiteboard eraser pick–wipe–place task under three demonstration settings: direct human demonstration, teleoperation without bilateral force feedback, and OmniUMI with bilateral gripper feedback and a handheld embodiment, measured by the $\ell_{2}$ norm of the 126-dimensional marker offset vector over time. OmniUMI produces tactile trajectories that are substantially closer to direct human demonstrations in both amplitude and temporal variation, whereas teleoperation yields noisier and less consistent tactile patterns.
As shown in Fig. 7, the tactile trajectory collected with OmniUMI more closely matches that of direct human demonstration, both in overall deformation magnitude and in temporal evolution. In contrast, the teleoperation baseline produces tactile signals that are more irregular and less consistent with the human reference, indicating reduced fidelity in contact perception and modulation during demonstration.
These results suggest that the human-aligned embodiment of OmniUMI improves not only internal grasp-force modulation and external-force regulation, but also tactile interaction acquisition. Without bilateral force feedback, teleoperation operators rely primarily on visual observation and therefore have greater difficulty reproducing the fine-grained temporal structure of human tactile interaction. By contrast, OmniUMI combines bilateral gripper feedback with a handheld embodiment, enabling more natural and physically grounded interaction with the environment and producing tactile trajectories whose temporal structure better matches human contact behavior.
Taken together, these results support the claim that human alignment should be evaluated across multiple physical interaction channels within the same contact-rich manipulation task rather than through vision alone. Across internal grasp force, external contact force, and tactile deformation, OmniUMI consistently produces demonstration trajectories that are closer to direct human behavior than conventional teleoperation, validating the effectiveness of its human-aligned multimodal interaction design.